a solution of 2z3+16i=0


1
Expert's answer
2022-05-31T16:23:33-0400
"2z^3+16i=0"

"z^3=-8i"

"z=2\sqrt[3]{-i}"

"-i=\cos(-\dfrac{\pi}{2})+i\sin(-\dfrac{\pi}{2})"

"k=0:"


"\sqrt[3]{1}(\cos(\dfrac{-\pi/2+2\pi(0)}{3})+i\sin(\dfrac{-\pi/2+2\pi(0)}{3}))"

"=\dfrac{\sqrt{3}}{2}-\dfrac{i}{2}"

"k=1:"


"\sqrt[3]{1}(\cos(\dfrac{-\pi/2+2\pi(1)}{3})+i\sin(\dfrac{-\pi/2+2\pi(1)}{3}))"

"=i"

"k=2:"


"\sqrt[3]{1}(\cos(\dfrac{-\pi/2+2\pi(2)}{3})+i\sin(\dfrac{-\pi/2+2\pi(2)}{3}))"

"=-\dfrac{\sqrt{3}}{2}-\dfrac{i}{2}"


"z_1=\sqrt{3}-i, z_2=2i,z_3=-\sqrt{3}-i,"